I greatly need your help. In trying to solve the problem. I found out that the hits are 25,000x24+110,000=710,000 hits. I'm having a hard time trying to set th prblem up as a proportions. Will you please help.
An administrator of a popular web site is told that a new faster server, which will replace the old server, can handle 45,000 "hits" (users accessing the site) per second. The web site currently experiences a peak demand of about 110,000 hits per second. But, every month, the peak demand increases by 2500 hits per second. Use a proportion equation to determine how many new servers the administrator should buy to address expected traffic for the next 24 months.

Answer by stanbon(48545)

An administrator of a popular web site is told that a new faster server, which will replace the old server, can handle 45,000 "hits" (users accessing the site) per second. The web site currently experiences a peak demand of about 110,000 hits per second. But, every month, the peak demand increases by 2500 hits per second. Use a proportion equation to determine how many new servers the administrator should buy to address expected traffic for the next 24 months.
-------------------------------------------
In 24 months the peak demand will be 110,000+2500*24 = 170,000
x/170,000 = 1/45000
x = 170/45
x = 3.777777
----------------
They will need 4 of the new servers.
Cheers,
Stan H.